OpenStudy (anonymous):

consider the vector I =[ 1/Square root (2) , 1/Square root (2)] and J= [ -1/Square root (2), 1/Square root (2)]. Write the vector (2,-6) in terms of I and J.

OpenStudy (phi):

james gave you a good starting point. I= [1 1]/sqrt(2) J= [-1 1]/sqrt(2) you want to find two scalars (two simple numbers) that you multiply I and J by to get (2,-6) call the numbers a and b then you must solve (2,-6)= a*I + b*J

OpenStudy (phi):

First rule: multiplying a vector by a scalar uses this rule: a (x,y)= (ax, ay)

OpenStudy (phi):

so, start with (2,-6)= a*I + b*J (2, -6) = a*(1,1)/sqrt(2) + b*(-1,1)/sqrt(2) multiply both sides by sqrt(2) . This means multiply every term by sqrt(2) you get sqrt(2) * (2, -6) = a*(1,1) + b*(-1,1)

OpenStudy (anonymous):

hmm this one doesn't make sense to me at all

OpenStudy (anonymous):

ohh ok

OpenStudy (phi):

Now use the rule First rule: multiplying a vector by a scalar uses this rule: a (x,y)= (ax, ay) on all the terms

OpenStudy (anonymous):

so its 2 sqrt(2) , -6 sqrt(2) = a (1,1) + b (-1,1)

OpenStudy (phi):

yes, but apply the rule to the two terms on the right hand side also

OpenStudy (anonymous):

how do you do that?

OpenStudy (phi):

a (x,y)= (ax, ay)

OpenStudy (anonymous):

so its 2sqrt(2) , -6sqrt(2) = (a1, a1) + (-b, b1)

OpenStudy (phi):

well, I would not write a1, because a*1 is just a.... also, put the left side in parens because we do not want to forget is is a vector

OpenStudy (anonymous):

ok. its [2sqrt(2), -6sqrt(2)] = (a,a) + (-b,b) ?

OpenStudy (phi):

yes, now add the two vectors on the right hand side

OpenStudy (anonymous):

how do you do that?

OpenStudy (phi):

You were able to subtract two vectors, for example P-Q in the last problem. Adding is the same, except you add instead of subtract.

OpenStudy (anonymous):

so its like (1+-1, 1+1) = (0,2) ?

OpenStudy (phi):

yes, but now we are adding variables instead of numbers

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

(a,a) + (-b,b) = ?? add corresponding entries

OpenStudy (anonymous):

(-ab, ab)

OpenStudy (phi):

add (not multiply)

OpenStudy (anonymous):

(a+-b, a+b)

OpenStudy (phi):

yes, but normally you would write a + -b as just a-b .

OpenStudy (anonymous):

oh ok. now what do i do?

OpenStudy (phi):

You have (2sqrt(2), -6sqrt(2)) = (a-b, a+b)

OpenStudy (phi):

this means you have two vectors that equal each other. If they equal each other, what does that mean about each entry in the vector?

OpenStudy (anonymous):

umm no clue

OpenStudy (phi):

if I told you ( 1, 2) = (x, y) what is x and what is y how do you know?

OpenStudy (anonymous):

that x = 1 and y =2

OpenStudy (phi):

Now use the same idea on our problem (2sqrt(2), -6sqrt(2)) = (a-b, a+b)

OpenStudy (anonymous):

that a-b = 2sqrt(2) and a+b = -6sqrt(2)

OpenStudy (phi):

yes. we have 2 equations and two unknowns. The way to solve these types of equations is write them down like this: a-b = 2sqrt(2) a+b = -6sqrt(2) add the two equations together (we are adding equal things to equal things, so the sum of them is still equal things) 2a-b+b= 2sqrt(2)-6sqrt(2) can you solve for a?

OpenStudy (anonymous):

lol thats confusing

OpenStudy (anonymous):

would it be a = 2sqrt(2) -6sqrt(2) / 2 ??

OpenStudy (phi):

first, you should put parens around (2sqrt(2) -6sqrt(2)) because you are dividing *both* terms by 2 (you divide the whole right hand side by 2) next, you should simplify

OpenStudy (anonymous):

so its [sqrt(2), -3sqrt(2)]?

OpenStudy (phi):

it is not a vector.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

oh sqrt(2)i , -3sqrt(2) j

OpenStudy (phi):

no. I mean you start with (2*sqrt(2)-6*sqrt(2))/2 which is just a (ugly) number: divide both terms by 2 to get (sqrt(2)-3*sqrt(2)) factor out sqrt(2) to get (1-3)*sqrt(2) simplify 1-3 to get a= -2*sqrt(2)

OpenStudy (phi):

btw, if you want info on how to solve simultaneous equations like this one, watch http://www.khanacademy.org/video/solving-systems-by-elimination?topic=worked-examples-4 and his other examples

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what would be the answer for this?

OpenStudy (phi):

we are getting there.... use a+b = -6sqrt(2) plus your value for a that we just found to get b

OpenStudy (anonymous):

so b = -6sqrt(2) - 2sqrt(2) ?

OpenStudy (phi):

well let's check that a + b= -6sqrt(2) -2*sqrt(2) + b= -6sqrt(2) add +2sqrt(2) to both sides to get b = -6sqrt(2)+2sqrt(2) so close, but not exactly correct. can you simplify this?

OpenStudy (anonymous):

lol i am so confused

OpenStudy (phi):

can you simplify b = -6sqrt(2)+2sqrt(2)

OpenStudy (anonymous):

umm -4sqrt(2) ?

OpenStudy (phi):

yes. we are almost done.

OpenStudy (phi):

we have to answer Write the vector (2,-6) in terms of I and J. which we decided meant: (2,-6)= a*I + b*J we now know a and b, so the final answer is?

OpenStudy (anonymous):

2sqrt(2) i + -4sqrt(2) j ?

OpenStudy (phi):

and what is I and J ? It does not hurt to write out the whole solution

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

I =[ 1/Square root (2) , 1/Square root (2)] J= [ -1/Square root (2), 1/Square root (2)] [2,-6]=2sqrt(2) I + -4sqrt(2) J

OpenStudy (phi):

You can check your answer by multiplying the scalars times the vectors, and then adding the vectors to see if you get [2, -6] btw, if you have time, you might want to brush up on your algebra using http://www.khanacademy.org/#core-algebra

OpenStudy (anonymous):

ohh k

OpenStudy (anonymous):

thanks a lot